OpenStudy (anonymous):

Solve 3x + 7 = -5 for x. Answer 4 -4 5 -5

OpenStudy (andriod09):

hey there. the first thing you would do is try to isolate x, which can be done by subtracting the 7 from both sides like so: \[3x+7=-5\]\[3x-7=-7-5\] and then the postive 7 and negative 7 next to the 3x would be canceled leaving you with \[3x=-12\] then what you would do is put them over 3, because you're now trying to isolate the x, \[\frac{3x}{3}=\frac{-12}{3}\] thus, making the answer -4, becuase when you divide a negative by a positive, the anwer is negative. Hope this helps!

OpenStudy (anonymous):

thanks!